40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 10 14:59:01 2008, rob, Update, Locking, RF common mode at zero offset lock_control_sequence_worked.pngtrendpowerbuild.png
    Reply  Fri Jul 11 23:55:25 2008, alan, Update, Locking, RF common mode at zero offset 
Message ID: 655     Entry time: Thu Jul 10 14:59:01 2008     Reply to this: 661
Author: rob 
Type: Update 
Category: Locking 
Subject: RF common mode at zero offset 
rob, john, yoichi

Last night we succeeded in reducing the CARM offset to zero.

We handed off control of the common mode servo from PO-DC to POX-I.

We pushed the common mode servo bandwidth to ~19kHz. Without the boosts, it had ~80 degs of phase margin. Didn't measure it after engaging the boosts (Boost + 1 superboost). Trying to engage the second superboost stage broke the lock.

The process is fully scripted, and the script worked all the way through several times.

The DARM ugf was ~200Hz. The RSE peak could clearly be seen. No optical spring, as expected (we're locking in anti-spring mode).

Engaging test mass de-whitening filters did not work (broke the lock).

I'm attaching a lock control sequence diagram and a trend of the arm power during a scripted up-sequence. I think the script can be sped up significantly (especially the long ramp period).

Up next:

Calibrated DARM spectrum
Noise hunting (start with dewhites)
DC - Readout
Lock to the springy side.
Attachment 1: lock_control_sequence_worked.png  85 kB  | Hide | Hide all
lock_control_sequence_worked.png
Attachment 2: trendpowerbuild.png  25 kB  | Hide | Hide all
trendpowerbuild.png
ELOG V3.1.3-